Nica's brother is now 3 times as old as she. In five years, his age will be twice hers. Find Nica's age now. Please provide an equation
Answer by ikleyn(52787)

.

Let x be the Nica's age.

Then the brother's age is 3x years.



In 5 years, the brother's age will be (3x+5) years,

while the Nica's age will be  x+5 years.



Therefore, the equation is

    3x + 5 = 2*(x+5).



If you want to solve it, simplify

    3x + 5 = 2x + 15

    3x - 2x = 15 - 5

       x    =   5


Thus Nica is 5 years old now.  The brother is 3*5 = 15 years old.

Solved, answered, explained.


Read attentively my solution line after line.


Enjoy learning (!)


---------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - HOW TO algebreze and to solve age problems?
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - Age problem for three participants
    - Miscellaneous age problems
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.